2) Full of hot air. A hot-air balloonist, rising vertically with a constant speed of 5.00 m/s, releases a sandbag at the instant the balloon is 60.0 m above the ground. After it is released, the sandbag encounters no appreciable air resistance a) Find the position and velocity of the sandbag at 0.200 s and 0.800 s after its release. (6) How much time does it take for the sandbag to strike the ground after it is released? (4) b)
Answer #1

Initial velocity of sandbag,110 = 5m/s

acceleration,Cl 9.81m/s

a)Position

Use Formula sut +1/2at2

y=u_{0}t-0.5gt^{2}

y5m/s*0.2s-0.5 * 9.81 * (0.2s)

y = 5 * 0.2-0.5 * 9.81 * 0.22

0.8038m

At 0.2s Position is 60.8038 m

=======================

Velocity

Use Formula v=u+at

v=u_{0}-gt

v -5m/s -9.81m/s2 0.2s

ANSWER: {color{Red} v_{0.2s}=3.038m/s}

======================================================================

Time =0.8s

Use Formula v=u+at

v=u_{0}-gt

Om/s 5m/s 9.81m/s2 t

t 0.51s

After t 0.51s sandbag is falling towards earth

----------------------------------

Use Formula sut +1/2at2

y=u_{0}t-0.5gt^{2}

y5m/s0.51s 0.59.81 (0.51s)2

y=5*0.51-0.5*9.81*0.51^{2}

y 1.2742 m

At t=0.51s Max height of 61.2742 m is reached

Time left =0.8-0.51s =0.29s

Position after falling for 0.29s

Use Formula sut +1/2at2

y=-0.5gt^{2}

y=-0.5*9.81*(0.29s)^{2}

y=-0.5*9.81*0.29^{2}

0.4125m

61.2742m-0.4125m=60.8617m

At 0.8s Position is 60.8617m

===================================================

Velocity

Use Formula v=u+at

v=u_{0}-gt

v=0m/s-9.81m/s^{2}*0.29s

ANSWER: {color{Red} v_{0.8s}=2.8449m/s}

====================================================

b)

Time to hit ground =0.51s +Time taken to fall from Max Height (61.2742m)

Use Formula sut +1/2at2

H_{max}=0-0.5gt^{2}

61.2742m = 0.5 * 9.81 *

t=12.49s

Time to hit ground =0.51s + 12.49s

ANSWER: {color{Red} t=13s}

================